# Reconfidencing LLMs from the Grouping Loss Perspective

### v2

@@ Abstract
Large Language Models (LLMs), such as GPT and LLaMA, are susceptible to generating hallucinated answers in a confident tone. While previous efforts to elicit and calibrate confidence scores have shown some success, they often overlook biases towards certain groups, such as specific nationalities. Existing calibration methods typically focus on average performance, failing to address this disparity. In our study, we demonstrate that the concept of grouping loss is an effective metric for understanding and correcting the heterogeneity in confidence levels. We introduce a novel evaluation dataset, derived from a knowledge base, specifically designed to assess the confidence scores of LLM responses across different groups. Our experimental results highlight significant variations in confidence, which are accurately captured by grouping loss. To tackle this issue, we propose a new method to calibrate the confidence scores of LLMs by considering different groups, a process we term reconfidencing. Our findings indicate that this approach effectively mitigates biases against minority groups, contributing to the development of fairer LLMs.
@@ 3.2 Experimental Settings
$P(\text{contradict}|s_{i},d)=\frac{\exp(z_{c})}{\exp(z_{e})+\exp(z_{c})}$ (1)
@@
$C_{\text{SelfCheckGPT}}(s_{i})=1-\frac{1}{m}\sum_{j=1}^{m}P(\text{contradict}|s_{i},d_{j})$ (2)
@@
Given the observed binary labels $Y$, the true posterior probabilities $Q$, confidence scores $C$ obtained from a model $P(Y)$, and the corresponding average true posterior probabilities $A$, the divergence of proper scoring rules can be decomposed as Kull and Flach (2015); Perez-Lebel et al. (2023):
$\mathbb{E}\left[f(S,Y)\right]=$ $\underbrace{\mathbb{E}\left[f(C,A)\right]}_{\text{Calibration Loss}}$ (3)
$+\underbrace{\mathbb{E}\left[f(A,Q)\right]}_{\text{Grouping Loss}}+\underbrace{\mathbb{E}\left[f(Q,Y)\right]}_{\text{Irreducible Loss}}$
where $f$ is a function that measures the divergence between the two inputs. In this work, we consider three metrics: the Brier Score $f^{\text{BS}}(S,Y)$ Brier (1950), the Calibration Loss $f^{\text{CL}}(S,C)$, and the Grouping Loss $f^{\text{GL}}(Q,Y)$  Kull and Flach (2015); Perez-Lebel et al. (2023). (1) The Brier score is the squared error between the observed binary labels $Y$ –denoting correct/incorrect answers– and the associated confidence scores $C$. The appealing property of the Brier score is that it is minimum when $C=P(y)$. (2) Calibration Loss (CL) measures the error rate (average observed $y$) for a given confidence score C: $\mathbb{E}[y|C=c]$; a calibration plot, as in Figure 2(a) plots this value for different values of $c$. When the confidence score $C$ equals the probability $P(y)$, the calibration plot is on the diagonal, and the calibration error is zero. However, the converse is not true: a calibration error can be zero and yet the confidence score differs from from the probability $P(y)$. The reason for this difference is that within the observations with a predicted confidence score of $C$, some have an actual probability above $C$ while others below: errors compensate Perez-Lebel et al. (2023). (3) Grouping Loss (GL) is the term missing to the calibration to fully control how the predicted confidence scores $C$ relate to the true probability $P(y)$. We reuse the method by Perez-Lebel et al. (2023) to estimate the lower bound of the grouping loss by looking at the dispersion in the error rate on sub-groups of observations for a given score $C$.
@@ 5 Conclusion
In this work, we analyzed how trustworthy current methods are when they give confidence scores to LLM answers. We create a novel dataset derived from the ground truth within the YAGO knowledge base, providing a framework for evaluating the calibration of confidence scores for different groups. Subsequent evaluations of different sizes of LLMs reveal a consistent discrimination towards particular minority groups. Leveraging estimators and visualizations, we show grouping loss in LLMs, such as those associated with long-tail entities and individuals of Asian origin. These findings emphasize that we should pay particular attention to minority groups when calibrating LLMs. Building upon these insights, we introduce a novel approach for reconfidencing LLMs based on latent sub-groups, resulting in improved calibrations. This new approach can mitigate the problem of hallucinations by generating alerts in response to LLM answers. Meanwhile, our findings can reduce biased information against groups such as race and gender, which is useful for the fairness of LLMs.

### v3

# Reconfidencing LLM Uncertainty from the Grouping Loss Perspective

@@ Abstract
[…]
@@ 3.2 Experimental Settings
$P(\text{contradict}|r_{i},d)=\frac{\exp(z_{c})}{\exp(z_{e})+\exp(z_{c})}$ (1)
@@
$S_{\text{SelfCheckGPT}}(r_{i})=1-\frac{1}{m}\sum_{j=1}^{m}P(\text{contradict}|r_{i},d_{j})$ (2)
@@
Given the observed binary labels $Y$, the true posterior probabilities $Q$, confidence scores $S$ obtained from a model $P(Y)$, and the corresponding average true posterior probabilities $C$, the divergence of proper scoring rules can be decomposed as Kull and Flach (2015); Perez-Lebel et al. (2023):
$\mathbb{E}\left[f(S,Y)\right]=$ $\underbrace{\mathbb{E}\left[f(S,C)\right]}_{\text{Calibration Loss}}$ (3)
$+\underbrace{\mathbb{E}\left[f(C,Q)\right]}_{\text{Grouping Loss}}+\underbrace{\mathbb{E}\left[f(Q,Y)\right]}_{\text{Irreducible Loss}}$
where $f$ is a function that measures the divergence between the two inputs. In this work, we consider three metrics: the Brier Score $f^{\text{BS}}(S,Y)$ Brier (1950), the Calibration Loss $f^{\text{CL}}(S,C)$, and the Grouping Loss $f^{\text{GL}}(C,Q)$  Kull and Flach (2015); Perez-Lebel et al. (2023). (1) The Brier score is the squared error between the observed binary labels $Y$ –denoting correct/incorrect answers– and the associated confidence scores $S$. The appealing property of the Brier score is that it is minimum when $S=P(y)$. (2) Calibration Loss (CL) measures the error rate (average observed $y$) for a given confidence score S: $\mathbb{E}[y|S=s]$; a calibration plot, as in Figure 2(a) plots this value for different values of $c$. When the confidence score $S$ equals the probability $P(y)$, the calibration plot is on the diagonal, and the calibration error is zero. However, the converse is not true: a calibration error can be zero and yet the confidence score differs from from the probability $P(y)$. The reason for this difference is that within the observations with a predicted confidence score of $S$, some have an actual probability above $C$ while others below: errors compensate Perez-Lebel et al. (2023). (3) Grouping Loss (GL) is the loss due to many instances being grouped under the same estimate $S$ while having different true posterior probabilities $Q$ Kull and Flach (2015). We reuse the method by Perez-Lebel et al. (2023) to estimate the lower bound of the grouping loss by looking at the dispersion in the error rate on sub-groups of observations for a given score $S$.
@@ 5 Conclusion
In this work, we analyzed how trustworthy current methods are when they give confidence scores to LLM answers. We create a novel dataset derived from the ground truth within the YAGO knowledge base, providing a framework for evaluating the calibration of confidence scores for different groups. Subsequent evaluations of different sizes of LLMs reveal a consistent discrimination towards particular minority groups. We show grouping loss in LLMs, such as those associated with long-tail entities and individuals of Asian origin. These findings emphasize that we should pay particular attention to minority groups when calibrating LLMs. Building upon these insights, we introduce a novel approach for reconfidencing LLMs based on latent sub-groups, resulting in improved calibrations. This new approach can mitigate the problem of hallucinations by generating alerts in response to LLM answers. Meanwhile, our findings can reduce biased information against groups such as race and gender, which is useful for the fairness of LLMs.
@@
Acknowledgements
This work was partially funded by projects NoRDF (ANR-20-CHIA-0012-01) and LearnI (ANR-20- CHIA-0026).